OpenStudy (anonymous):

e^(x/10) = -y^2 + 100y how would I make y the subject

OpenStudy (anonymous):

add y^2 to both sides

OpenStudy (anonymous):

does that mean "solve for y'?

OpenStudy (anonymous):

yes, y = ..

OpenStudy (anonymous):

then subtract e^(x/10) to both sides

OpenStudy (dumbcow):

you could treat it like a quadratic where e^x/10 is your "c" term \[y= 50\pm \sqrt{2500 - e ^{x/10}}\]

OpenStudy (anonymous):

y^2 = -e^(x/10) + 100y

OpenStudy (anonymous):

and this leads to dumbcows answer,,lol

OpenStudy (anonymous):

0 = -y^2 + 100y -e^(x/10) quadratic form of equation

OpenStudy (anonymous):

y^2 -100y +e^(x/10) =0

OpenStudy (anonymous):

y=(-(-100)-+sqrt(100^2 -4(1)(4e^(x/10))))/2 y=(100-+sqrt(10000-4e^(x/10)))/2 y=(50-+sqrt(2500-e^(x/10))